OpenStudy (anonymous):

Factor: 2x^2+5x+3 I need to learn the steps.

jhonyy9 (jhonyy9):

-5+/- 1 x_1,_2= --------- = -3/2 and -1 4

jhonyy9 (jhonyy9):

us quadratic formula

OpenStudy (anonymous):

so far i have 2x^2 +2x+3x+3 now what do i do?

jhonyy9 (jhonyy9):

2x(x+1) +3(x+1) =(x+1)(2x+3)

jhonyy9 (jhonyy9):

t

jhonyy9 (jhonyy9):

this is with factorised method

jhonyy9 (jhonyy9):

is right too

jhonyy9 (jhonyy9):

but you can using the quadratic formula too -b+/- sqrt(b2 x_1,_2= ------------

OpenStudy (anonymous):

okay i know the answer is (2x+3)(x+1) idk how to do these

jhonyy9 (jhonyy9):

2x+3=0 2x=-3 x=-3/2

jhonyy9 (jhonyy9):

and x+1=0 x=-1

OpenStudy (anonymous):

okay. Do you know if you can help me with this one? 3n^2+7n+4

OpenStudy (anonymous):

I just have to factor

OpenStudy (anonymous):

try this 2x^2 +5x+3=0 x^2 +(5/2)x+(3/2)=0 (x+ 3/2)(x +1)=0 x=-3/2 and x=-1

OpenStudy (anonymous):

directly from factoring

jhonyy9 (jhonyy9):

write 7n in the form 3n+4n

OpenStudy (anonymous):

i did

jhonyy9 (jhonyy9):

and this is ok ?

OpenStudy (anonymous):

lets see if i do this: 3n(n+1) +4(n+1) =(n+1)(3n+4) Will that give me: (3n+4) (n+1)?

OpenStudy (anonymous):

3n^2+7n+4=0 n^2+(7/3)n+4/3=0 (n+ 4/3)(n+1)=0 n=-4/3 and x=-1

jhonyy9 (jhonyy9):

yes is right

OpenStudy (anonymous):

correct also ..lol